A Bibliography of Electronic Publishing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 January 2019 Version 2.06

Abstract

This bibliography records books and articles on electronic publishing. It includes papers from the proceedings of several recent conferences on the topic. ISBN and ISSN numbers are given, where available.

Coverage of the proceedings of EP’94 and RIDT’94 can be found in a separate bibliography, epodd.bib, covering the journal Electronic Publishing—Origination, Dissemination, and Design.

Title word cross-reference

12th [Ano96e]. 1991 [MA91]. 1992 [SOFS93]. 1993 [Shi93]. 1996 [Ano96m].

20/20 [Sec97]. 2nd [Ano93e].

4th [HJB98].

5th [Ano97b].

6th [Ano94b].

7-10 [VC92a]. 7th [HJB98].

86th [Ano96j]. ’89 [Ano89].

’92 [VC92a]. ’96 [Ano96a, Ano96f]. ’98 [HJB98].

AAS [Ano93f]. AAUP [Bar95, Cre95]. abcdefg [Ada89]. Abstract [Roh86]. Academic [Per95]. Access [Baj97, Glu97, Lyo95]. Accuracy [Sen93]. achieve [MR92]. Acrobat [SBEH93]. Active [EJM^{+}90]. Activities [Ano93j]. ADAPT [HW90]. Adaptive [ZB93]. administrator [McN93]. Adobe [Har93].

Advanced [ABH93, HB91]. After [Rub93]. Aftershocks [Bur95]. Age [Ano93h, Spe93]. ahead [Shi93]. Aids [Pit95]. Aleph [SW92]. Alerting [SOFS93]. Algorithm [CP89, IO93]. Algorithms [OO89]. aliased [AMS89]. Aliasing [NN89]. Alternative [Har93]. Alto [BBKF95]. AMS [KOOR93]. Amsterdam [Ano97b]. Analysing [HH93]. Analysis [BAC93, HW90, LB95, WSG89]. Annual [Ano93e, Ano96j]. Anti [AMS89, NN89]. Anti-aliased [AMS89]. Anti-Aliasing [NN89]. Anything [Jen97]. Anyway [Luc97]. Appearance [UHST89]. Application [BH89, Cla89]. Applications [Bro95, MNBH92, VES95, LG97]. Approach [DL91, FX93, Gon91, HI93, NNGP89, Ver90, Won86]. Approaches [CX97, vv91]. April [AGS93, Ano93g, Ano96e, HJB98, HMQ93, Shi93, VC92a, van88]. Arabic [Har95a]. Architecture [FSD92]. Archive [Sea93]. Area [Cla89]. Art [Die94]. artistic [HJB98]. ASHTEX [Gal86]. ASIS [LG97]. Assessment [AS89]. Associates [Ano93d]. Association [Ano96j]. Astronomical [Van93]. Astronomy [Ano93f]. Australia [Ano93g]. Author [GG97]. authors [BKDH86]. Auto [Gon91]. Auto-tracing [Gon91]. Automated [MR92]. Automatic [AYO91, Aho95, And90, HW90, Kar89, SAB92]. avant [Leb92]. avant-texte [Leb92]. Awarded [Ano93j]. aware [Bro95].

Based [ERE92, FSD92, NN89, SB95, SOFS93, VM95, ZD95, BLR92, FW93, MBM93, WSG89]. be [Hof95]. Berkeley [Pit95]. Better [Ada89]. between [DEB93, HS93]. Beyond [Ano97a, Ano97b]. Bézier [CP89]. BILETA [Ano96a]. Binary [BB92]. Bitmapped [Maa89]. Bitmaps [BB92, Ros91]. Block [SB95]. Block-Based [SB95]. Book [SS97]. Bookstore [Bar95]. Boston [MA91]. Boundary [Ano93o]. Breaking [Ham97, HG88]. Brink [Ano93c]. Broadsheet [LB92]. Brother [Ano93a].

Browsing [SOFS93]. Brush [UHST89].

Brush-written [UHST89].

CA [BBKF95]. CAJUN [SBEH93]. Calligraphic [Doo89]. CaminoReal [ABMW88]. Campus [Ano93b, Sea95]. Can [Ano93a, MR92]. Canberra [Shi93]. capabilities [PH86]. case [Hay93]. Casual [Ano93a]. Catalog [Bar95, Cre95]. Catalog/Bookstore [Bar95]. Cataloging [Jen97]. Cause [Ano93b]. CD [Ano94b]. CD-ROM [Ano94b]. Center [Sea93]. Centralization [Lam97]. Centre [Ano90]. Century [Ano93h, Har95b]. Change [Ano96b, Cox93]. Changes [Shi93]. Changing [Ano93h]. Character [AP91, Fan91, HH93, IO93, Sou91]. Characters [AMS89, Gon89, Maa89, OO89, UHST89]. Chinese [MS90, CX97, DL91, Fan91, UHST89]. Chip [CP91]. Choosing [BB92]. Chosen [Ano93d]. Client [FSD92]. Client-Server-Based [FSD92]. Cluster [Coo91]. Collect [BLR92]. Collect-and-Stitch [BLR92]. Colophon [AM91, AS93, Fur90b, MH93, VC92b]. Color [Ami91, Coo91]. Columbia [Ano93m]. Combined [Bar95]. commercial [Ano85]. Communication [SH93, Shi93]. Communications [Lyo95, McM95, Sec97, Cox93]. Community [SH93]. Comparative [Baj97]. Compilation [GB93]. Complementarity [Ver90]. Complex [Sou92]. Complexity [Ros91]. Composition [Sou92, Zhe92]. Compressing [Ros91]. Compression [ZD95]. Computer [ERE92, SOFS93]. Computer-Based [ERE92]. Computers [Mid97]. Comserve [SH93]. Concept [Law95]. Concepts [ABH93]. Concise [Ano93m]. Conference [AGS93, Ano93c, Ano93g, Ano93o, Ano94b, BBKF95, Dés86, Fur90a, HJB98, HMQ93, SM96, VC92a, van88, Ano85, Ano86b, Ano86a, Ano87, Ano89, Ano90, Ano93e, Ano97b, Shi93, Ano96j, Ano97a]. Conferences [AH89]. consensus [Ano96l]. Consequences [BB92]. Constraint [Ada89, MBM93]. Constraint-based [MBM93]. Constraints [BH89, Erf93, HB91, Sen93, Ano86a]. Construction [MUP91]. Containment [CP89]. Content [Aho95, SAB92, TC92]. Context [BWB92, Bro95]. Context-aware [Bro95]. Context-Dependent [BWB92]. Continuous [Cla89]. Continuous-tone [Cla89]. Contour [Fah89]. contributed [LG97]. Control [Mac91, SB93]. Conversion [Duf89]. Converting [Aer95, Zhe91]. Convolution [Duf89]. Cooperating [Cha90]. Cooperation [Ver90]. Coordinate [Due93]. Coordinate-independent [Due93]. copy [BPKL88]. Copyright [GG97, Hof95, Oak93, Sam97, Sou95]. Corporate [Ano88, Ano85]. Correction [RFP92]. Cost [Key95, Lam93]. Council [Sec97]. Coventry [Ano96a]. Craft [MR92]. Creating [Bro95]. Creation [CJ90, RMF93]. Cross [Ano93o]. Cross-Boundary [Ano93o]. CRT [Nai91]. Cuadra [Ano93d]. Cubics [Gon91]. Cultivation [Ele95]. Cultural [Bus97]. Current [Die94]. Curve [IO93]. Curves [CP89]. Customized [CJ90].

Darmstadt [AGS93, HMQ93]. Database [ABH93, Ano93d, KLMN90, RH90]. Databases [Ano94b]. Day [Rub93]. DC [Oke93, Oke95]. Dead [Arn93a]. Debate [Lam97]. December [Ano86a, Oke93]. Decentralization [Lam97]. Decipherability [Tam91]. Declarative [NNGP89]. Dedicated [CP91]. Defined [Doo89]. Delivery [CJ90]. demand [McL93]. Dependent [BWB92]. describe [LT86]. Description [Due93, Sou91].

Descriptive [CJ90]. Design [Bar95, BH93, DL91, Fan91, KLLP90, Koh89, NNGP89]. Designing [Sou86]. Detection [AP91]. Determination [SAB92]. Developing [MNBH92]. Dialogue [Ano93o]. Dictionary [Aer95]. Differences [DEB93]. Different [vv91]. Diffusion [ZB93]. Digital [AA89, AH89, AGS93, Coo91, Lyo95, MA91, SB93, Ano96c, Ano96g, Ano96k, HJB98]. Digitalisation [Tam89]. Digitization [AP91, MS90]. directives [LT86]. Disability [Ano93l]. Disciplinary [EI95]. Discrete [CP89]. Discrimination [Dys93]. Displacement [AYO91]. Display [AMS89, Dys93, LB92]. Distributed [FSD92, dGH89]. Distribution [Ano94a, Van93]. DMBS [LG90]. Do [GG97, Mid97]. Document [Arn93b, BB90, BWB92, BBKF95, Bro95, FX93, FS88, Fur90a, HW90, HI93, HMQ93, Ker90, KLMN90, Lev93, Lev95, LB95, PR92, QNA90, Spi90, VC92a, ZD95, Zhe92, van88, BKDH86, CCH^{+}86b, Fer86a, Pet86]. Documentation [Dés86, TC92]. Documents [BG90, BCO92, CJ90, EJM^{+}90, Erf93, FW93, FL93, GB93, HS93, Kin95, KP93, KS95, LG90, Mun95, MH92, RA95, RV93, RH90, BPKL88, CFB86, LT86]. Does [But95]. Domain [WSG89]. Domains [KFA95]. Dots [Coo91]. dpi [Mac91]. Draft [Rod95]. Driven [Ada89]. Dynamic [AB89, Sta93a].

e-MED [Aer95]. E-prints [Kel95]. EasyTEX [CFB86]. Ecole [AH89]. economics [Ano86a, McL93]. Economy [Bra95]. EDDS [PH86]. Edge [AP91]. edited [Sec97]. Editing [AQ92, KB90, Ver90]. Editor [BLR92, MBM93, RFP92, WW95]. Editors [Cha90]. Education [Sol93]. Efficient [Fah89, HI93, Koh89]. Electrical [Sec97]. Electronic [Aer95, Ano85, Ano86b, Ano86a, Ano90, Ano93e, Ano93c, Ano93f, Ano93g, Ano93h, Ano93k, Ano93m, Ano94b, Ano96a, Ano96e, Ano96d, Ano97a, Baj97, BEG90, BBKF95, But95, But94, CX97, CJ90, Day95, Die94, Ele95, Esp93, Fur90a, Glu97, Gue95, Hay93, HJB98, HMQ93, Jen97, KOOR93, Lam97, LB92, LG97, MGLD97, Oak93, Owe93, RMF93, Rou93, Rub93, Sea93, Sea95, Sec97, SM96, SWB97, SS97, SH93, SOFS93, Van93, VC92a, van88, Ano87, Ano88, Ano89, Ano96l, Ano97b, Cox93, McN93, Oke93, Spe93, Ano96k, Sea93]. Elsevier [Die94]. Encoding [Hoc93]. Encyclopedia [Ano93m]. Engineering [QNA90, Sec97]. Engineers [Sec97]. English [Aer95]. Environment [Ada89, KS95, Lev95, Zhe92, dGH89, App86, CCH^{+}86a, CCH^{+}86b, Pet86, QVB86]. EP [Ano87, Ano88, HJB98]. EP89 [Ano89]. EP90 [Ano90]. EP92 [VC92a]. Error [ZB93]. especially [Mac91]. European [Dés86]. Evaluation [Baj97, MUP91]. Evaluations [Tam89]. Evolutional [Tam89]. Evolving [Ano93a]. Ex [Ano94b]. Exact [Duf89]. Examines [Ano93b]. Example [Due93, MGLD97]. Executive [Ano93i, Sou95]. Exhibition [Ano90]. Experience [Bil89, BEG90]. Experiences [FSD92, Pet86]. Experiment [KOOR93, Van93]. Expert [SM96]. Extended [Har95a]. Extensibility [Sta93b]. Extensible [EJM^{+}90].

Facet [All95]. factory [Pet86]. Fair [Har95b, Law95, Sou95]. Family [Ano93b]. FAQs [VM95]. Fast [AMS89, Gon89, Gon91]. February [SM96]. Federal [VC92a, Ano93h]. Fédérale [AH89]. Fifth [HMQ93]. Files [LB95]. Filled [Gon89]. Filling [Fah89, Oke95]. Filtering [KS95]. Finding [Pit95]. First [Har95b]. FirstSearch [Ano93m]. Fiscal [Ano93j]. fit [LB92]. Fitting [IO93]. Five [Rou93, Sei95]. Five-Year [Rou93]. Flexible [BH89]. flo [WB90]. Flowcharts [WB90]. Font [AYO91, Ada89, AS89, BH93, CP91, DL91, Due93, Her89, Kar89, Kla93, MB93, NN89, NNGP89, OO89, Ros91, Sta93b, ZI93, dGH89]. Font-scaler [Sta93b]. Fonts [And90, AB89, BB89, Har93, HP93, IO93, Koh89, Mac91, MS90, NN89, Sta93a, Tam89]. Formats [Sea95]. Formatter [MH92, WW95]. Formatting [Sou92, Van92, LT86]. formulae [CFB86, Luc86]. Forum [SGBOP93, Sec97]. Foundation [BB90]. fourth [Oke95]. Framework [Bro95, MNBH92]. France [Dés86, HJB98, SM96, van88]. Frankenstein [Ele95]. Frequency [WSG89]. Full [TR92]. Full-Text [TR92]. funding [McN93]. Future [Ano94a, Die94].

Gaithersburg [Fur90a]. Galileo [Bur95]. Generation [AYO91, Ada89, Aho95, And90, Gon89, MB93, OO89, Oke93, Won86]. Geographic [Glu97]. Germany [AGS93, HMQ93]. Get [But95]. Glyphs [Har95a]. Good [Mid97]. Gopher [Ano93l]. Government [Sou95]. Grammar [BB90, FW93, KP93]. Grammar-based [FW93]. Grants [Ano93j]. Graphics [BLR92, DL91]. Gray [AYO91]. Gray-Scale [AYO91]. Grayscale [Far91]. Grayscale/Resolution [Far91]. Greyscale [BB92]. Grid [BH89, HB91, WSG89]. Grid-based [WSG89]. Gridfitting [And90]. GRIF [QVB86, RFP92]. Group [Sou95].

Halftone [Coo91, MUP91]. Halftoning [Coo91, SB93]. Handling [ERE92]. Hands [BCR93]. Hands-on [BCR93]. Hard [BPKL88]. Hebrew [Tam89, Tam91]. HEIRAlliance [Ano93i]. held [Ano85, Ano86b, Ano86a, Ano87, Ano88, Ano89, HJB98, MA91, Sec97]. Help [Sam97]. Helvetica [DEB93]. Heritage [Bus97]. Hexagonal [WSG89]. Hexagonal- [WSG89]. Hierarchy [MH92]. High [AMS89, AYO91, Cla89, PH86]. High-Quality [Cla89, AMS89]. Highest [MR92]. Hinder [Sam97]. Hinting [Kar89]. Hints [And90]. HortBase [MGLD97]. Hotel [Oke93, Oke95]. HTML [Sea95]. Hypermedia [KLLP90]. Hypertext [FSD92, HS93, Ver90]. HyTime [Erf93, FFE95]. hz [Zap93]. hz-program [Zap93].

i.e [Ano94b]. ICSU [SM96]. Identity [Lev95]. IEEE [SOFS93]. If [Jen97]. II [MA91]. Image [Cla89, Far91, SOFS93]. Image-Based [SOFS93]. Images [HI93, KB90, WSG89]. Imaging [AH89, AGS93, MA91, HJB98]. Immunity [BB89]. impact [Cox93]. implementation [Pet86]. Implications [LB95, SS97, LG97]. imprint [Ano96g]. improved [CCH^{+}86a]. Improving [Dys93, Sen93]. Incremental [FX93, NNGP89]. independent [Cla89, Due93]. industry [Ano93e]. Infinifont [MB93]. Information [Ano93b, Ano94a, Baj97, Bra95, Bus97, Glu97, MGLD97, Ano96i, McL93]. Initiative [Hoc93]. Innovation [Key95, Ano96b]. input [CFB86]. INRSTEX [Fer86a]. Institute [VC92a]. Institution [Sec97]. Integrating [PH86]. Integration [LG90, Ano87]. Intellectual [Sou95, Spe93]. Intelligent [Fan91, Kar89, Sta93a]. Intensity [Zhe91]. Interactive [CJ90, KFA95, SGBOP93, App86, ABMW88, CFB86, Gal86, QVB86]. Interest [Sou95]. Interface [HS93]. Interfaces [BG90, KFA95]. International [AH89, AGS93, Ano94b, Ano96e, Ano96f, BBKF95, Fur90a, HJB98, HMQ93, VC92a, van88, Ano85, Ano86b, Ano87, Ano96l]. Internet [Ano94a, Ano96f, BCR93, CX97, GG97]. Introduction [Her89, O’D93]. Investigation [Maa89]. invited [BWB92, Leb92, Van92]. ISI [Ano96k].

Issues [Hof95, KLLP90, Ker90].

Japanese [AMS89]. Job [Ano93b]. John [Ano93l]. Joint [SM96]. jointly [HJB98]. Journal [KOOR93, Sei95, SBEH93, Sol93, SGBOP93]. Journals [Gue95, Kel95, LK95, Ano96e]. June [Ano87, Ano88, Ano89, Ano90, Ano94b, Dés86].

Kanji [AYO91, Due93, OO89, Ros91]. Karlsruhe [BKDH86]. KEYNOTE [Per95]. know [Ano96h]. Knowledge [KFA95]. Knowledge-rich [KFA95].

Labyrinth [EI95]. Language

[Arn93b, Mun95, SAB92, SW92, WB90, Zhe92, CJ90, LT86]. languages [Fer86a]. Laser [Mac91]. Lausanne [AH89, VC92a]. Law [Lyo95]. Lawmakers [Hof95]. Layout [BB90, PR92]. learned [Ano96e]. Learning [NNGP89]. Legibility [BB89, Tam91]. level [BAC93]. L’hypertexte [Leb92]. Librarians [Ano93h]. Libraries [Ano96j, Fri97, Gue95, Mac97, McM95, Sea93]. Library [Ano93j, Ano96k, SOFS93, Ano93g, Ano96c, McN93]. Limitations [HB91]. line [HG88, Sea93]. linearities [Nai91]. Linearization [Nai91]. Linguistic [MBM93]. Lita [Ano93k]. Literate [Hof95]. Literature [Van93]. Live [Arn93a]. Lively [vv91]. Logic [BB90]. Logical [BAC93, HI93, RV93]. London [Ano85, Ano86b, Ano86a, Ano87, Ano88, Ano89, Ano90, Ano96e]. Long [Arn93a]. Look [Ano93a]. Looking [Mac91]. Lossy [ZD95]. Low [BAC93, Dys93, NN89, Ros91]. Low-level [BAC93]. Lucida [BB89]. Luminance [Nai91].

Malo [HJB98]. Management [dGH89]. Managing [Cha90, Cox93]. Manipulated [MUP91]. Manipulation [BBKF95, FL93, Fur90a, HMQ93, VC92a, van88].

Manufacture [Sou91]. March [Ano96a, HJB98]. March/April [HJB98]. MarkItUp [FX93]. Markup [CJ90, Roh86]. marvellous [Gal86]. Maryland [Fur90a]. Master [Har93]. Matching [HH93]. mathematical [ABMW88, Luc86]. May [LG97, Sec97]. MED [Aer95]. Mediating [HS93]. Medieval [EI95]. meeting [LG97, MA91]. Merging [RV93]. METAFONT [Har93, Sou86]. Metaphor [ERE92]. Method [Fah89, NN89, Ros91, TR92, UHST89, Zhe91]. Methods [LB95]. s [Mac91]. Micro [Zap93]. Micro-typography [Zap93]. Middle [Koh89, Aer95]. Middlesex [Ano90]. midyear [LG97]. Minneapolis [LG97]. Mixed [HI93]. MN [LG97]. model [Pet86]. Modelling [Kin95, TC92]. Models [Aho95, RH90]. Modern [Tam91]. Moiré [Ami91]. Monograph [Arn93a]. Multi [Coo91, Zhe92]. Multi-Cluster [Coo91]. Multi-Language [Zhe92]. Multilingual [RFP92, Spi90, Fer86b]. Multimedia [CJ90, Erf93, Hof95, Kin95, Ano93e, SGBOP93]. Multiple [Har93, dGH89, Fer86a]. Muse [LK95]. Music [FL93, HG88].

Namable [Lev95]. Names [Lev95]. Naming [Lev95]. Natural [SAB92]. Natural-Language [SAB92]. Navigation [Pit95]. Need [Har95b, Hof95]. Network [RMF93]. Networked [Lev95]. Networking [Ano93j]. networks [Oke93]. News [AA89, LB92, SWB97]. Newsletter [Ano93k]. Next [But94]. Nice [van88]. niche [Ano96h]. NII [Sou95]. Noise [BB89]. Non [Nai91]. Non-linearities [Nai91]. Normalizing [UHST89]. not-for-profit [Oke93]. notebook [ABMW88]. November [Ano85, Oke95]. NY [Ano93e].

Object [BCO92, EJM^{+}90, FFE95, RMF93, Sta93b].

Object-orientation [Sta93b]. Object-Oriented [BCO92, EJM^{+}90]. Objects [Lyo95]. OCLC [Ano93m]. Oct [Ano97b, MA91]. October [AH89, Ano93e]. ODA [BPKL88, GG90]. ODA-documents [BPKL88]. Offering [Ano93m]. On-line [Sea93]. One [Sei95]. Online [Bar95, Cre95, SGBOP93]. opportunities [Oke93]. Optical [ZI93]. Optimal [HG88]. organised [Sec97]. Organization [Ele95]. orientation [Sta93b]. Oriented [BCO92, EJM^{+}90]. Outcomes [Shi93]. Outline [And90, BH89, CP91, Gon89, Sta93a]. Outlined [CP89]. Outlines [HP93, vv91]. overheads [Esp93].

Page [GG90]. Pagination [BKKW95]. Palo [BBKF95]. Pandora [Bil89]. Paper [Ano93b, ERE92, Mid97, BWB92, Leb92, Van92]. Papers [MA91, LG97]. Paradigm [Ano93a]. Paradigms [Ano93a, FFE95]. Parametric [DL91, Gon91, MB93]. Parametrization [Har93]. Paris [SM96]. part [Sec97]. Past [Ano93a, O’D93]. Patent [Hof95]. Pattern [ZD95]. Pattern-Based [ZD95]. patterns [MUP91, Shi93]. paying [Oke95]. Pays [Bro97]. PDF [LB95, SB95]. Peer [Luc97]. Pen [Doo89]. Performance [DEB93]. Performances [HB91]. periodicals [Ano96i]. permission [SOFS93]. Perseus [Myl93]. Perspective [Die94, Lyo95, McN93]. PHAROS [CP91]. Phenomenon [Ami91]. Physical [RV93]. pipeline [Oke95]. piper [Oke95]. pixel [AP91]. Pixels [Mac91]. Plan [Rou93]. Plane [CP89]. Planner [GG90]. Plans [Ano93f]. please [BKDH86]. Point [CP89]. Polygon [Duf89]. Polytechnique [AH89]. Possible [Bar97, Kla93]. PostScript [Har93]. Power [MUP91]. Practice [BEG90]. Preparation [Ker90, BKDH86, CCH^{+}86b, Fer86a]. Preprints [Ano94b, Kel95]. Present [Ano93a, Ano94a, O’D93]. Presentation [Mun95, Sou92]. Preserving [Ano96g]. Press [Ano93n, SM96]. Press/Unesco [SM96]. Presses [Gue95]. previewer [Gal86]. Pricing [Day95, VM95]. Printed [Ano97a, Ano97b]. Printer [Mac91]. Printing [Cla89]. prints [Kel95]. Problems [AS89]. Proceedings [AH89, AGS93, BBKF95, Dés86, Fur90a, HMQ93, VC92a, van88, Ano85, Ano86b, Ano86a, Ano87, Ano88, Ano89, Ano90, HJB98, Oke93, Oke95, Sec97, SM96]. process [Pet86]. Processing [BWB92, HW90, Spi90, Zhe92, App86]. Product [Die94]. Production [Ano93d]. Products [Day95, Hay93]. Professional [GG90, MGLD97]. Profiles [ZI93]. profit [Oke93]. program [Zap93]. Programme [Sec97]. Progress [Cre95]. Project [Ano93d, Bar95, Cre95, McM95, Sei95, Ano93n, Ano96k, Bur95, LK95, Myl93, Pit95, SBEH93]. Projection [ZI93]. Projects [Ano93f]. Properties [Cha90]. Property [Sou95, Spe93]. Proposal [Rod95]. Prospects [ABH93, Ano86a]. Provision [McL93]. Publication [Law95]. Publications [KLLP90]. Publisher [Ano96h]. Publishers [McM95, Owe93, BKDH86]. Publishing [ABH93, Ano93c, Ano93f, Ano94b, Ano97a, Bra95, BBKF95, But95, But94, CX97, Esp93, Fur90a, GG90, HMQ93, Kel95, KOOR93, Lam97, Oak93, Per95, Rod95, Rou93, Sea95, SBEH93, SS97, VC92a, van88, Ano85, Ano86b, Ano86a, Ano87, Ano88, Ano89, Ano90, Ano93e, Ano93g, Ano96a, Ano96e, Ano96d, Ano96l, Ano97b, HJB98, LG97, Oke93, Sec97, SM96, BEG90]. Pursuit [MR92].

Qualitative [BAC93]. Quality [AYO91, AS89, Cla89, Far91, Mac91, MR92, AMS89]. Quasi [Doo89]. Quasi-Calligraphic [Doo89]. Queries [Arn93b].

Raster

[AH89, AGS93, HJB98, MA91, ZB93].

Rasterization [And90, Her89]. Rasterizing [HP93]. Rational [CP89]. Readability [Tam91]. Reading [DEB93]. Recognition [FX93, HI93, Sen93, Spi90, ZI93]. Reconsidered [BKKW95]. Reconstruction [BWB92, PR92]. Recovering [PR92]. Recovery [Key95]. Red [Ano93n, But94]. Redux [Ele95]. Reference [RMF93]. Regions [CP89]. Regularisation [Sta93a]. Relationships [SAB92]. Release [Ano93n]. Releases [Ano93i]. remainder [Zhe91]. Renaissance [Fri97, Mac97]. Rendering [CP91]. Rendition [Doo89, BPKL88]. Report [Ano93i, Cre95, Sou95]. Reports [Ano93c]. Repositories [FFE95]. Representation [BB90, Cla89, FL93, Kla93]. Representing [RMF93]. reprinted [SOFS93]. Research [But95]. Resolution [Cla89, Dys93, Far91, Koh89, NN89, Zhe91, PH86]. Resolution-independent [Cla89]. Resource [Ano93j]. Resource-Sharing [Ano93j]. response [Ano93g]. Results [LB95]. Retrieval [All95, TR92]. Retrieving [Ano96i, Luc86]. Reuse [Lev93]. Review [Luc97]. Reviews [SS97]. Reward [But95]. rich [KFA95]. Riding [Bur95]. RIDT [HJB98, MA91]. RightPages [SOFS93]. rights [Ano93e, Ano96b, Spe93, Sou95]. Role [CJ90]. Roles [Ano93h, MGLD97]. ROM [Ano94b]. Rules [Sou92]. Running [App86].

Sage [Ano93n, But94]. Scale [AYO91]. scaler [Sta93b]. Scaling [Kar89]. Scan [Duf89]. ScEX [FL93]. Scheme [ZD95]. Scholarly [Arn93a, Bra95, But95, McM95, Oke93, Owe93, Rod95, Ano93g, McL93, Shi93]. Scholars [MBM93, SS97, SH93]. Scholarship [But95, Ele95]. School [Per95]. science [Ano96d, Ano96l, SM96, Die94].

Sciences [Kel95]. Scientific [Dés86, KOOR93, Rod95, Ano96i, CFB86]. Scrimshaw [Arn93b]. Script [Doo89]. Second [Ano93i, Dés86, MA91, Oke93]. seminar [Ano96e]. Sensible [Bar97]. Separate [GB93]. Separation [Ami91]. September [Ano86b, BBKF95, Fur90a, SOFS93]. Server [Bar95, EI95, FSD92]. services [Hay93]. Session [BCR93]. Set [Har95a]. Setting [Lam97]. SGML [Aho95, BWB92, Bus97, FFE95, Hoc93, LT86, MNBH92, VES95]. SGML/HyTime [FFE95]. Shape [Maa89, Noo91]. Shapes [HH93]. Sharing [Ano93j]. Sheets [ERE92, RA95]. Silicon [CP91]. SilverPlatter [Ano93d]. SIMON [FW93]. Simple [Koh89]. Sixth [BBKF95]. Size [UHST89]. Societie [MGLD97]. Societies [Sei95]. Software [Ano93d]. Solutions [KLLP90]. Some [Tam89, VM95, BB92]. Spaces [All95]. Sparks [Ano93o]. Spatial [ERE92, Glu97, Nai91, Zhe91]. Special [Ano96j]. Specification [Erf93]. Specifying [FS88]. Spectra [MUP91]. Spline [Fah89]. Splines [Kla93]. Spreading [Ano96k]. Square [WSG89]. Square- [WSG89]. SSP [Ano93o]. St [HJB98, Ano93h]. St. [Ano93l]. Standardized [Sea95]. Standards [Glu97, Hoc93, MR92, Pit95]. Starting [Sol93]. State [Die94]. Statistics [Sol93]. STELAR [Van93]. Step [But94]. steps [Pet86]. Stick [Bro95]. Stick-e [Bro95]. Still [Mid97]. Stitch [BLR92]. Strasbourg [Dés86]. Strategies [Ano93i, Mid97]. Strategy [Die94, Esp93]. Strengths [RH90]. String [HH93, RFP92]. String-to-String [RFP92]. Stroke [AYO91, Noo91]. Structure [FX93, HI93, NN89, PR92, RMF93]. Structured [AQ92, BCO92, FW93, FS88, FSD92, GB93, HS93, KLMN90, KP93, KS95, LG90, Mun95, MH92, RFP92, RA95, SB95, Van92, Ver90].

Structures [BAC93, RV93]. Studies [EI95]. study [Hay93]. Sub [AP91]. Sub-pixel [AP91]. Subdivision [Zhe91]. Subdivision-remainder [Zhe91]. Summary [BCR93, Sou95, KOOR93]. supply [McL93]. Support [Ano93j]. Supporting [CP91]. Swiss [VC92a]. Switzerland [AH89, VC92a]. Sydney [Ano93g]. Symbols [Dys93]. Symposium [Dav95, O’D93, Oke93, Oke95]. SYNDOC [KS95]. Synopsis [Dav95]. System [But95, Cha90, EJM^{+}90, FW93, FSD92, KLMN90, MB93, PR92, dGH89, BKDH86, Fer86a]. Systems [AQ92, Ker90, Lev93, MGLD97, Sen93].

Tables [Van92]. Tabular [WW95].

Tackling [LK95]. Tagging [HW90]. Task [DEB93]. Technical [CX97, TC92]. Techniques [HH93, Sou91]. Technologies [dGH89]. Technology [Ano93j, VC92a]. TEI [Sea95]. Telecommunications [Ano93a]. Temporal [Erf93]. Text [ERE92, KB90, Owe93, Sea93, Sen93, Sou92, TR92, App86, Hoc93]. texte [Leb92]. Texts [Rub93, SAB92, Sea93]. Textual [RH90]. Texture [SB93]. Their [MS90]. theme [Sec97]. theological [Won86]. thinking [Lam93]. third [AGS93, Won86]. Time [O’D93]. Times [DEB93]. Tokyo [Ano94b]. Tomorrow [Rub93]. tone [Cla89]. Toolkit [BCO92]. Topography [MA91]. tracing [Gon91]. Tradeoffs [Far91, Ker90]. Tradition [Ham97]. Traditional [Har95a]. Trajectory [Doo89]. Transformation [AQ92, FW93, KP93]. Transformations [Arn93b, FS88]. Trends [AA89]. Tulip [But94]. Turmoil [Fri97, Mac97]. Tutorial [BCR93]. Twenty [Har95b]. Twenty-First [Har95b]. Type [AQ92, Sou91]. Typecase [Har95a]. typeface [Sou86]. Typefaces [Tam91]. Typesetting [MR92, SW92, WB90, Won86]. Typographic [RA95]. Typographical [Sen93]. Typography [AA89, AH89, AGS93, BBKF95, Fur90a, HMQ93, MR92, VC92a, van88, HJB98, Zap93].

U.S. [Sou95]. UK [Ano86a, Ano90, Ano96a]. Unesco [SM96]. Unfilled [Gon89]. Unicode [BH93, Har95a]. University [Gue95, Lam97, Pet86, Ano93l, Sea93]. USA [BBKF95, Oke93]. Usage [VM95]. Usage-Based [VM95]. Use [Har95b, KP93, Law95, Sou95, SS97]. User [BG90, CCH^{+}86a, Mid97]. Users [BB92]. Using [AYO91, Coo91, Due93, ABH93, Erf93, Sen93].

Variable [Kla93]. VAT [Ano96m]. Verge [Ano93c]. Verheissung [Won86]. versatile [BKDH86]. Version [Aer95, Ano93k]. Versions [Lev95]. Versprechen [Won86]. very [PH86]. Virginia [Sea93]. Virtual [Law95]. Vision [KOOR93, Sec97]. visions [Oke93]. Vista [Oke93, Oke95]. VORTEX [CCH^{+}86b]. vs [Lam97].

Wants [Luc97]. Washington [Oke93, Oke95]. Way [Koh89, Shi93]. Weaknesses [RH90]. Weaving [RMF93]. Web [RMF93, EI95, VES95]. Wembley [Ano90]. Where [But95]. Who [Bro97, Luc97]. Whom [Bro97]. Wide [EI95]. Width [Kla93]. Will [Har95b, Sam97, Sec97]. Word [Ano97a, Ano96k, Ano97b]. Work [RMF93, Sec97]. Working [Sou95]. World [Ano96f, EI95, Jen97, Ano97b, Gal86, Ano97a]. Writing [ERE92]. written [UHST89].

Xtable [WW95].

Year [Ano93j, Rou93]. Yeti [BLR92]. York [Ano93e].

References

Adams:1989:NTD

[AA89] D. Adams and J. André. News trends in digital typography. In André and Hersch [AH89], pages 14–21. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Andre:1989:DF

[AB89] J. André and B. Borghi. Dynamic fonts. In André and Hersch [AH89], pages 198–204. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Aberer:1993:PPU

[ABH93] K. Aberer, K. Boehm, and C. Hueser. The prospects of publishing using advanced database concepts. In Hüser et al. [HMQ93], pages 469–480. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Arnon:1988:CIM

[ABMW88] Dennis Arnon, Richard Beach, Kevin McIsaac, and Carl Waldspurger. CaminoReal: an interactive mathematical notebook. In van Vliet [van88], pages 1– 18. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Adams:1989:ABC

[Ada89] D. Adams. abcdefg: A better constraint driven environment for font generation. In André and Hersch [AH89], pages 54– 70. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Aertsen:1995:TEC

[Aer95] H. Aertsen. Towards an e-MED:

Converting the Middle English dictionary into an electronic version. In Okerson [Oke95], pages 125–134. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Andre:1993:PTI

[AGS93] Jacques André, Jakob Gonczarowski, and Richard Southall, editors. Proceedings of the third International Conference on Raster Imaging and Digital Typography, 11–13 April 1994, Darmstadt, Germany, volume 6(3) of Electronic Publishing—Origination, Dissemination, and Design. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Andre:1989:RID

[AH89] Jacques André and Roger D. Hersch, editors. Raster Imaging and Digital Typography: Proceedings of the International Conferences, Ecole Polytechnique Fédérale, Lausanne, Switzerland, October 1989. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Ahonen:1995:AGS

[Aho95] H. Ahonen. Automatic generation of SGML content models. In Brown et al. [BBKF95], pages 195–206. ISSN 0894-3982.

Allen:1995:RFS

[All95] R. B. Allen. Retrieval from facet spaces. In Brown et al. [BBKF95], pages 247–258. ISSN 0894-3982.

Andre:1991:C

[AM91] Jacques André and Robert A. Morris. Colophon. In Morris and André [MA91], pages 216– 217. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Amidror:1991:MPC

[Ami91] Isaac Amidror. The moiré phenomenon in color separation. In Morris and André [MA91], pages 98–120. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Abe:1989:FDH

[AMS89] S. Abe, M. Maekawa, and K. Shimuzu. Fast display of highquality anti-aliased Japanese characters. In André and Hersch [AH89], pages 134–143. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Andler:1990:AGG

[And90] Sten F. Andler. Automatic generation of gridfitting hints for rasterization of outline fonts. In Furuta [Fur90a], pages 221–234. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Anonymous:1985:EPC

[Ano85] Anonymous, editor. Electronic publishing: corporate & commercial publishing: proceedings of the international conference held in London, November 1985. Online Publications, London, UK, 1985. ISBN 0-86353-039-7. LCCN Z286.E43E436 1985. UK£55.00.

Anonymous:1986:EPU

[Ano86a] Anonymous, editor. Electronic publishing & the UK: prospects, economics & constraints: proceedings of a conference held in London in December 1984. British Library, London, UK, 1986. ISBN 0-7123-3063-1.

Anonymous:1986:EPP

[Ano86b] Anonymous, editor. Electronic publishing: proceedings of the international conference held in London, September 1986, volume 2 of Electronic media series. Online Publications, London, UK, 1986. ISBN 0-86353-058-3. LCCN Z286.E43E439 1986.

Anonymous:1987:EEP

[Ano87] Anonymous, editor. EP: electronic publishing: towards integration: proceedings of the international conference held in London, June 1987, volume 9. Online Publications, London, UK, 1987. ISBN 0-86353-093-1. LCCN Z286.E43E474 1987.

Anonymous:1988:CEP

[Ano88] Anonymous, editor. Corporate electronic publishing: proceedings of EP held in London, June 1988. Online Publications, London, UK, 1988. ISBN 0-86353-151-2. LCCN Z286.E43E6 1988.

Anonymous:1989:EEP

[Ano89] Anonymous, editor. EP89: electronic publishing ’89: proceedings of the conference held in London, June 1989. Blenheim Online Ltd., London, UK, 1989. ISBN 0-86353-176-8. LCCN Z286.E43E433 1989.

Anonymous:1990:EPC

[Ano90] Anonymous, editor. Electronic publishing conference proceedings: EP90, 19–21 June, 1990, Wembley Exhibition and Conference Centre, London, Middlesex, UK. Blenheim Online Ltd., London, UK, 1990. ISBN 0-86353-211-X. LCCN Z 286 E43 E6 1990.

Anonymous:1993:BCY

[Ano93a] Anonymous. “brother, can you paradigm?: A casual look at past, present, and evolving telecommunications paradigms”. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 22–?? ISSN 0741-9058.

Anonymous:1993:CPE

[Ano93b] Anonymous. Cause paper examines information job family on campus. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 30–?? ISSN 0741-9058.

Anonymous:1993:CRE

[Ano93c] Anonymous. Conference reports: Electronic publishing “on the Verge of a Brink”. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 12–14. ISSN 0741-9058.

Anonymous:1993:CAS

[Ano93d] Anonymous. Cuadra Associates software chosen by SilverPlatter for new database production project. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 31–?? ISSN 0741-9058.

Anonymous:1993:API

[Ano93e] Anonymous, editor. Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY, number 108 in Library Hi Tech News. Pierian Press, Ann Arbor, MI, USA, December 1993. ISSN 0741-9058.

Anonymous:1993:EPA

[Ano93f] Anonymous. Electronic publishing in astronomy; projects and plans of the AAS. In Okerson [Oke93], pages 133–144. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Anonymous:1993:ESP

[Ano93g] Anonymous, editor. Electronic scholarly publishing: the library response: Conference — April 1993, Sydney, Australia, volume 42(2) of Australian Library Journal. Australian Library and Information Association, Sydney, NSW, Australia, 1993. ISSN 0004-9670.

Anonymous:1993:FLS

[Ano93h] Anonymous. Federal librarians in the 21 st century: Changing roles in the electronic age. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 16–?? ISSN 0741-9058.

Anonymous:1993:HRS

[Ano93i] Anonymous. HEIRAlliance releases second executive strategies report. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 19–?? ISSN 0741-9058.

Anonymous:1993:LTG

[Ano93j] Anonymous. Library technology grants awarded to support networking and resource-sharing activities for fiscal year 1993. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 1–?? ISSN 0741-9058.

Anonymous:1993:LNE

[Ano93k] Anonymous. Lita Newsletter in electronic version. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 31–?? ISSN 0741-9058.

Anonymous:1993:NDG

[Ano93l] Anonymous. New disability gopher at St. John’s University. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 27–?? ISSN 0741-9058.

Anonymous:1993:ONO

[Ano93m] Anonymous. OCLC now offering Concise Columbia Electronic Encyclopedia on FirstSearch. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 30–?? ISSN 0741-9058.

Anonymous:1993:PRR

[Ano93n] Anonymous. Press release on Red Sage Project. In Okerson [Oke93], pages 119–122. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Anonymous:1993:SCS

[Ano93o] Anonymous. SSP conference sparks cross-boundary dialogue. In Electronic publishing and rights: 2nd Annual publishing industry conference on electronic publishing and multimedia — October 1993, New York, NY [Ano93e], pages 14–?? ISSN 0741-9058.

Anonymous:1994:IDT

[Ano94a] Anonymous. Information distribution through Internet: Its present and future. In Preprints / Electronic Publishing Conference. The 6th International Conference on Electronic Conference [i.e., Publishing, Ex The International Conference on CD-ROM Databases], June 17, 1994, Tokyo [Ano94b], pages 26–40. ISBN 4-89700-005-X. LCCN ????

Anonymous:1994:PEP

[Ano94b] Anonymous, editor. Preprints / Electronic Publishing Conference. The 6th International Conference on Electronic Conference [i.e., Publishing, Ex The International Conference on CD-ROM Databases], June 17, 1994, Tokyo, Electronic Publishing Conference -preprints-. USACO Corporation, Tokyo, Japan, 1994. ISBN 4-89700-005-X. LCCN ????

Anonymous:1996:BEP

[Ano96a] Anonymous, editor. BILETA ’96: Electronic publishing, March 1996, Coventry, UK. British and Irish Legal Education Technology Association, CTI Law Technology Centre, University of Warwick, Coventry, UK, 1996.

Anonymous:1996:CIR

[Ano96b] Anonymous. Change and innovation in rights. In Shaw and Moore [SM96], pages 245–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:DL

[Ano96c] Anonymous. The digital library.

In Shaw and Moore [SM96], pages 225–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:EPS

[Ano96d] Anonymous. Electronic publishing in science. In Shaw and Moore [SM96], pages 233–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:EPI

[Ano96e] Anonymous, editor. Electronic publishing: International learned journals seminar; 12th — April 1996, London, volume 9(4). Association of Learned and Professional Society Publishers, Hatfield, Herts, UK, 1996. ISSN 0953-1513.

Anonymous:1996:IWI

[Ano96f] Anonymous. Internet World International 96. In Shaw and Moore [SM96], pages 237–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:PDI

[Ano96g] Anonymous. Preserving the digital imprint. In Shaw and Moore [SM96], pages 215–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:PKY

[Ano96h] Anonymous. Publisher, know your niche! In Shaw and Moore [SM96], pages 203–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:RIS

[Ano96i] Anonymous. Retrieving information from scientific periodicals. In Shaw and Moore [SM96], pages 219–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:SLA

[Ano96j] Anonymous. Special Libraries Association 86th Annual Conference. In Shaw and Moore [SM96], pages 242–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:SDW

[Ano96k] Anonymous. Spreading the digital word: ISI’s Electronic Library Project. In Shaw and Moore [SM96], pages 211–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:TIC

[Ano96l] Anonymous. Towards an international consensus on electronic publishing in science. In Shaw and Moore [SM96], pages 207–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1996:V

[Ano96m] Anonymous. VAT 1996. In Shaw and Moore [SM96], pages 243–?? ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Anonymous:1997:BPWa

[Ano97a] Anonymous. Beyond the printed word: The World Electronic Publishing Conference. In Beyond the printed word: the world electronic publishing conference: 5th — 1997 Oct: Amsterdam [Ano97b], pages 22–33.

Anonymous:1997:BPWb

[Ano97b] Anonymous, editor. Beyond the printed word: the world electronic publishing conference: 5th — 1997 Oct: Amsterdam. IFRA, ????, 1997.

Avrahami:1991:SPE

[AP91] Gideon Avrahami and Vaughn Pratt. Sub-pixel edge detection in character digitization. In Morris and André [MA91], pages 54– 64. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Appelt:1986:RTI

[App86] W. Appelt. Running TEX in an interactive text processing environment. In Désarménien [Dés86], pages 1–8. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Akpotsui:1992:TTS

[AQ92] Extase Akpotsui and Vincent Quint. Type transformation in structured editing systems. In Vanoirbeek and Coray [VC92a], pages 27–41. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Arnold:1993:SMD

[Arn93a] K. Arnold. The scholarly monograph is dead; long live the scholarly monograph. In Okerson [Oke93], pages 73–80. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Arnon:1993:SLD

[Arn93b] D. S. Arnon. Scrimshaw: A language for document queries and transformations. In Hüser et al. [HMQ93], pages 385–396. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Adams:1989:PFQ

[AS89] D. Adams and R. Southall. Problems of font quality assessment. In André and Hersch [AH89], pages 213–222. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Andre:1993:C

[AS93] J. Andre and R. Southall. Colophon. In André et al. [AGS93], pages 307–308. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Abe:1990:HQG

[AYO91] Hiroshi Abe, Yoshimichi Yamamoto, and Yoshio Ohno. High quality gray-scale Kanji font generation using automatic stroke displacement. In Morris and André [MA91], pages 147–155. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Belaid:1993:QAL

[BAC93] A. Belaid, J. C. Anigbogu, and Y. Chenevoy. Qualitative analysis of low-level logical structures. In Hüser et al. [HMQ93], pages 435– 446. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Bajjaly:1997:AEI

[Baj97] S. T. Bajjaly. Access to electronic information: A comparative evaluation. In Logan and Gluck [LG97], pages 107–112. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Barton:1995:CAO

[Bar95] B. H. Barton. The combined AAUP online catalog/bookstore project: Server design. In Okerson [Oke95], pages 87–88. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Barron:1997:PS

[Bar97] D. Barron. Possible, but sensible? In Secker [Sec97], pages 15– 22. ISBN 0-85296-691-1. LCCN ????

Bowden:1989:NIL

[BB89] P. R. Bowden and D. F. Brailsford. On the noise immunity and legibility of Lucida fonts. In André and Hersch [AH89], pages 205–212. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Brown:1990:LGF

[BB90] Allen L. Brown, Jr. and Howard A.

Blair. A logic grammar foundation for document representation and document layout. In Furuta [Fur90a], pages 47–64. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Black:1992:CBG

[BB92] Alison Black and Andrew Boag.

Choosing binary or greyscale bitmaps: some consequences for users. In Vanoirbeek and Coray [VC92a], pages 247–260. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Brown:1995:PSI

[BBKF95] Allen Brown, Anne Brüggemann-Klein, and An Feng, editors. Proceedings of the Sixth International Conference on Electronic Publishing, Document Manipulation and Typography, 24–26 September, 1996, Palo Alto, CA, USA, volume 8(2) of Electronic Publishing—Origination, Dissemination, and Design. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995. ISSN 0894-3982.

Brown:1992:OOT

[BCO92] Heather Brown, Fred Cole, and Elizabeth Oxborrow. An objectoriented toolkit for structured documents. In Vanoirbeek and Coray [VC92a], pages 95–111. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Burton:1993:HIT

[BCR93] D. R. Burton, K. W. Curnow, and D. L. Rodgers. Hands-on Internet tutorial session; A summary. In Okerson [Oke93], pages 15– 18. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Brailsford:1990:EPP

[BEG90] David F. Brailsford, David R. Evans, and Geeti Granger. Electronic Publishing—practice and experience. In Furuta [Fur90a], pages 169–182. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Bier:1990:DUI

[BG90] Eric A. Bier and Aaron Goodisman. Documents as user interfaces. In Furuta [Fur90a], pages 249–262. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Betrisey:1989:FAO

[BH89] C. Bétrisey and R. D. Hersch. Flexible application of outline grid constraints. In André and Hersch [AH89], pages 242– 250. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Bigelow:1993:DUF

[BH93] C. Bigelow and K. Holmes. The design of a Unicode font. In André et al. [AGS93], pages 289– 306. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Billawala:1989:PEM

[Bil89] N. Billawala. Pandora: An experience with . In André and Hersch [AH89], pages 34– 53. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Bruggemann-Klein:1986:HPA

[BKDH86] A. Bruggemann-Klein, P. Dolland, and A. Heinz. How to please authors and publishers: a versatile document preparation system at Karlsruhe. In Désarménien [Dés86], pages 9–31. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Bruggemann-Klein:1995:PR

[BKKW95] A. Brüggemann-Klein, R. Klein, and S. Wohlfeil. Pagination reconsidered. In Brown et al. [BBKF95], pages 139–152. ISSN 0894-3982.

Brotsky:1992:YGE

[BLR92] Daniel Brotsky, David M. Levy, and Dan Rabin. Yeti: a graphics editor based on “collect-and-stitch”. In Vanoirbeek and Coray [VC92a], pages 211–221. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Behrmann-Poitiers:1988:HCR

[BPKL88] J. Behrmann-Poitiers, H. Keil, and H. Loebl. Hard copy rendition of ODA-documents. In van Vliet [van88], pages 71–82. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Braman:1995:SPI

[Bra95] S. Braman. Scholarly publishing in the information economy. In Okerson [Oke95], pages 39– 50. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Brown:1995:SDF

[Bro95] P. J. Brown. The Stick-e document: A framework for creating context-aware applications. In Brown et al. [BBKF95], pages 259–272. ISSN 0894-3982.

Brown:1997:WPW

[Bro97] D. Brown. Who pays what to whom. In Secker [Sec97], pages 41–50. ISBN 0-85296-691-1. LCCN ????

Burr:1995:RAG

[Bur95] E. Burr. Riding the aftershocks:

The Galileo Project. In Okerson [Oke95], pages 119–124. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Busch:1997:SCH

[Bus97] J. A. Busch. SGML for cultural heritage information. In Logan and Gluck [LG97], pages 9– 26. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Butter:1994:NSE

[But94] K. Butter. The next step in electronic publishing: Red Sage and Tulip. In Anonymous [Ano94b], pages 1–15. ISBN 4-89700-005-X. LCCN ????

Butler:1995:RRS

[But95] J. Butler. Research into the reward system of scholarship; where does scholarly electronic publishing get you? In Okerson [Oke95], pages 167–178. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Brown:1992:RCD

[BWB92] Allen L. Brown, Jr., Toshiro Wakayama, and Howard A. Blair. A reconstruction of contextdependent document processing in SGML (invited paper). In Vanoirbeek and Coray [VC92a], pages 1–25. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Chen:1986:IUE

[CCH^{+}86a] P. Chen, J. Coker, M. A. Harrison, J. W. McCarrell, and S. Procter. An improved user environment for TEX. In Désarménien [Dés86], pages 32– 44. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Chen:1986:VDP

[CCH^{+}86b] P. Chen, J. Coker, M. A. Harrison, J. W. McCarrell, and S. Procter. The VORTEX document preparation environment. In Désarménien [Dés86], pages 45–54. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Crisanti:1986:ETI

[CFB86] E. Crisanti, A. Formigoni, and P. La Bruna. EasyTEX: Towards interactive formulae input for scientific documents input with TEX. In Désarménien [Dés86], pages 55–64. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Chamberlin:1990:MPS

[Cha90] Donald D. Chamberlin. Managing properties in a system of cooperating editors. In Furuta [Fur90a], pages 31–46. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Cruz:1990:RDM

[CJ90] Gil C. Cruz and Thomas H. Judd. The role of a descriptive markup language in the creation of interactive multimedia documents for customized electronic delivery. In Furuta [Fur90a], pages 277–290. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Clauer:1989:RIA

[Cla89] A. A. Clauer. A resolutionindependent, area continuoustone image representation and its application for high-quality printing. In André and Hersch [AH89], pages 271–286. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Cook:1991:CDH

[Coo91] Craig C. Cook. Color digital halftoning using multi-cluster halftone dots. In Morris and André [MA91], pages 120–127. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Cox:1993:IEC

[Cox93] J. Cox. The impact of electronic communications: Managing change. In Anonymous [Ano93g], pages 104–107. ISSN 0004-9670.

Corthout:1989:PCA

[CP89] M. Corthout and E.-J. Pol. A point containment algorithm for regions in the discrete plane outlined by rational Bézier curves. In André and Hersch [AH89], pages 169–180. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Corthout:1991:SOF

[CP91] Marc Corthout and Evert J.

Pol. Supporting outline font rendering in dedicated silicon: the PHAROS chip. In Morris and André [MA91], pages 177– 189. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Creesy:1995:AOC

[Cre95] C. Creesy. The AAUP online catalog project: A progress report. In Okerson [Oke95], pages 89– 92. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Chen:1997:CEP

[CX97] H. Chen and H. Xu. Chinese electronic publishing on the Internet: Technical approaches. In Logan and Gluck [LG97], pages 55– 66. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Davis:1995:SS

[Dav95] J. Davis. A synopsis of the symposium. In Okerson [Oke95], pages 1–14. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Day:1995:PEP

[Day95] C. Day. Pricing electronic products. In Okerson [Oke95], pages 51–56. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

DeLange:1993:PDB

[DEB93] R. W. De Lange, H. L. Esterhuizen, and D. Beatty. Performance differences between Times and Helvetica in a reading task. In André et al. [AGS93], pages 241–248. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Desarmenien:1986:TSD

[Dés86] Jacques Désarménien, editor. TEX for Scientific Documentation. Second European Conference. Strasbourg, France, June 19–21, 1986. Proceedings, number 236 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

deBry:1989:MMF

[dGH89] R. deBry, A. W. Griffee, and J. P.

Hofmeister. Management of multiple font technologies in a distributed system environment. In André and Hersch [AH89], pages 223–231. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Dietz:1994:ESE

[Die94] R. D. J. Dietz. Elsevier Science electronic product strategy: Current state of the art and future perspective. In Anonymous [Ano94b], pages 17–24. ISBN 4-89700-005-X. LCCN ????

Dong:1991:PGA

[DL91] Yunmei Dong and Kaide Li. A parametric graphics approach to Chinese font design. In Morris and André [MA91], pages 156– 165. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Dooijes:1989:RQC

[Doo89] E. H. Dooijes. Rendition of quasicalligraphic script defined by pen trajectory. In André and Hersch [AH89], pages 251–260. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Duerst:1993:CFD

[Due93]  M. J. Duerst.  Coordinate-

independent font description using Kanji as an example. In André et al. [AGS93], pages 133– 144. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Duff:1989:PSC

[Duf89] T. Duff. Polygon scan conversion by exact convolution. In André and Hersch [AH89], pages 154– 168. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Dyson:1993:IDS

[Dys93] M. C. Dyson. Improving discrimination of symbols for display at low resolution. In André et al. [AGS93], pages 231–240. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Everhart:1995:LWW

[EI95] D. Everhart and M. Irvine. The Labyrinth: A World Wide Web disciplinary server for medieval studies. In Okerson [Oke95], pages 33–38. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

English:1990:EOO

[EJM^{+}90] Paul M. English, Ethan Jacobson, Robert A. Morris, Kimbo B. Mundy, Stephen D. Pelletier, Thomas A. Polucci, and H. David Scarbro. An extensible, objectoriented system for active documents. In Furuta [Fur90a], pages 263–276. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Eleey:1995:FRO

[Ele95] M. Eleey. Frankenstein redux: Organization and cultivation of electronic scholarship. In Okerson [Oke95], pages 27–32. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Eklundh:1992:WSP

[ERE92] Kerstin S. Eklundh, Staffan Romberger, and Per Englund. Writing on sheets of paper: a spatial metaphor for computer-based text handling. In Vanoirbeek and Coray [VC92a], pages 143–151. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Erfle:1993:STC

[Erf93] R. Erfle. Specification of temporal constraints in multimedia documents using HyTime. In Hüser et al. [HMQ93], pages 397–412. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Esposito:1993:TSE

[Esp93] J. J. Esposito. Towards a strategy for electronic publishing (overheads). In Okerson [Oke93], pages 107–118. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Fahlander:1989:SCM

[Fah89] O. Fahlander. A spline contour method with efficient filling. In André and Hersch [AH89], pages 83–96. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Fan:1991:TIC

[Fan91] Jianping Fan. Towards intelligent Chinese character design. In Morris and André [MA91], pages 166– 176. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Farrell:1991:GRT

[Far91] Joyce E. Farrell. Grayscale/ resolution tradeoffs in image quality. In Morris and André [MA91], pages 65–80. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Ferguson:1986:IDP

[Fer86a]  M. J. Ferguson.  INRSTEX:

a document preparation system for multiple languages. In Désarménien [Dés86], pages 74– 88. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Ferguson:1986:MT

[Fer86b] M. J. Ferguson. A multilingual TEX. In Désarménien [Dés86], pages 65–73. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Francois:1995:SHR

[FFE95] P. François, P. Futtersack, and C. Espert. SGML/HyTime repositories and object paradigms. In Brown et al. [BBKF95], pages 63– 80. ISSN 0894-3982.

Filgueiras:1993:RMM

[FL93] M. Filgueiras and J. P. Leal. Representation and manipulation of music documents in ScEX. In Hüser et al. [HMQ93], pages 507– 518. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Friend:1997:LTR

[Fri97] F. Friend. Libraries: Turmoil and renaissance. In Secker [Sec97], pages 59–64. ISBN 0-85296-691-1. LCCN ????

Furuta:1988:SSD

[FS88] Richard Furuta and P. D. Stotts.

Specifying structured document transformations. In van Vliet [van88], pages 109–120. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Furuta:1992:ECS

[FSD92] Richard Furuta, P. David Stotts, and Gregory D. Drew. Experiences with a client-server-based architecture for a distributed structured hypertext system. In Vanoirbeek and Coray [VC92a], pages 113–125. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Furuta:1990:PIC

[Fur90a] R. Furuta, editor. Proceedings of the International Conference on Electronic Publishing, Document Manipulation & Typography, Gaithersburg, Maryland, September 1990. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Furuta:1990:C

[Fur90b] Richard Furuta. Colophon. In Furuta [Fur90a], pages 291–293. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Feng:1993:SGT

[FW93] A. Feng and T. Wakayama. SI-MON: A grammar-based transformation system for structured documents. In Hüser et al.

[HMQ93], pages 361–372. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Fankhauser:1993:MIA

[FX93] P. Fankhauser and Y. Xu. Mark-ItUp! an incremental approach to document structure recognition. In Hüser et al. [HMQ93], pages 447–456. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Gallot:1986:AIP

[Gal86] L. Gallot. ASHTEX: An interactive previewer for TEX or the marvellous world of ASHTEX. In Désarménien [Dés86], pages 89– 97. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Groves:1993:SCS

[GB93] M. J. Groves and D. F. Brailsford. Separate compilation of structured documents. In Hüser et al. [HMQ93], pages 315–326. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Guardalben:1990:OPP

[GG90] Giovanni Guardalben and Mosé Giacomello. An ODA page planner for professional publishing. In Furuta [Fur90a], pages 79–92. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Gregory:1997:CIW

[GG97] V. L. Gregory and W. S. Gregory. Copyright on the Internet: What’s an author to do? In Logan and Gluck [LG97], pages 131– 141. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Gluck:1997:SEA

[Glu97] M. Gluck. Standards for electronic access to geographic and spatial information. In Logan and Gluck [LG97], pages 27– 40. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Gonczarowski:1989:FGU

[Gon89] J. Gonczarowski. Fast generation of unfilled and filled outline characters. In André and Hersch [AH89], pages 97–110. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Gonczarowski:1991:FAA

[Gon91] Jakob Gonczarowski. A fast approach to auto-tracing (with parametric cubics). In Morris and André [MA91], pages 1–15. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Guedon:1995:EJL

[Gue95] J.-C. Guedon. Electronic journals, libraries, and university presses. In Okerson [Oke95], pages 67–76. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Hammond:1997:BT

[Ham97] R. Hammond. Breaking with tradition. In Secker [Sec97], pages 9– 14. ISBN 0-85296-691-1. LCCN ????

Haralambous:1993:PPF

[Har93] Y. Haralambous. Parametrization of PostScript fonts through METAFONT: an alternative to Adobe Multiple Master fonts. In André et al. [AGS93], pages 145– 158. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Haralambous:1995:TAT

[Har95a] Y. Haralambous. The traditional Arabic typecase extended to the Unicode set of glyphs. In Brown et al. [BBKF95], pages 111–124. ISSN 0894-3982.

Harper:1995:WWN

[Har95b] G. Harper. Will we need fair use in the Twenty-First Century? In Okerson [Oke95], pages 195– 212. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Hayes:1993:ESP

[Hay93] H. Hayes. Electronic services and products: A case study. In Anonymous [Ano93g], pages 89– 93. ISSN 0004-9670.

Hersch:1991:AGC

[HB91] Roger D. Hersch and Claude Bétrisey. Advanced grid constraints: Performances and limitations. In Morris and André [MA91], pages 190–204. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Hersch:1989:IFR

[Her89] R. Hersch. Introduction to font rasterization. In André and Hersch [AH89], pages 1–13. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Hegazy:1988:OLB

[HG88] Wael A. Hegazy and John S. Gourlay. Optimal line breaking in music. In van Vliet [van88], pages 157–169. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Herz:1993:ACS

[HH93] J. Herz and R. D. Hersch.

Analysing character shapes by string matching techniques. In André et al. [AGS93], pages 261– 272. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Hu:1993:MAT

[HI93] T. Hu and R. Ingold. A mixed approach toward an efficient logical structure recognition from document images. In Hüser et al. [HMQ93], pages 457–468. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Hersch:1998:EPA

[HJB98] Roger Hersch, André Jacques, and Heather Brown, editors. Electronic publishing, artistic imaging, and digital typography: 7th International Conference on Electronic Publishing, EP ’98 held jointly with the 4th International Conference on Raster Imaging and Digital Typography, RIDT ’98, St. Malo, France, March/April 1998: proceedings, volume 1375 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN ???? ISSN 0302-4743. LCCN ????

Huser:1993:PFI

[HMQ93] Christoph Hüser, Wiebke Möhr, and Vincent Quint, editors. Proceedings of the Fifth International Conference on Electronic Publishing, Document Manipulation and Typography, 13–15 April, 1994, Darmstadt, Germany, volume 6(4) of Electronic Publishing—Origination, Dissemination, and Design. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1993. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Hockey:1993:ESS

[Hoc93] S. Hockey. Encoding standards: SGML and the Text Encoding Initiative: What and why? In Okerson [Oke93], pages 59– 64. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Hofstetter:1995:MPC

[Hof95] F. T. Hofstetter. Multimedia patent and copyright issues: The need for lawmakers to be multimedia literate. In Okerson [Oke95], pages 183–188. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Hussain:1993:ROF

[HP93] F. Hussain and M. L. V. Pitteway. Rasterizing the outlines of fonts. In André et al. [AGS93], pages 171–182. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Hayashi:1993:MIB

[HS93] K. Hayashi and A. Sekijima. Mediating interface between hypertext and structured documents. In Hüser et al. [HMQ93], pages 423–434. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Handley:1990:AAD

[HW90] John Handley and Stuart Wood.

ADAPT: Automatic document analysis processing and tagging. In Furuta [Fur90a], pages 183– 192. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Itoh:1993:CFA

[IO93] K. Itoh and Y. Ohno. A curve fitting algorithm for character fonts. In André et al. [AGS93], pages 195–206. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Jeng:1997:WIA

[Jen97] Ling Hwey Jeng. What, if anything, is cataloging in an electronic world? In Logan and Gluck [LG97], pages 41– 54. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Karow:1989:AHI

[Kar89] P. Karow. Automatic hinting for intelligent font scaling. In André and Hersch [AH89], pages 232– 241. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Kopec:1990:EIT

[KB90] Gary E. Kopec and Steven C.

Bagley. Editing images of text. In Furuta [Fur90a], pages 207–220. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Kelly:1995:PEP

[Kel95] B. Kelly. Publishing E-prints, preprints, and journals in the sciences. In Okerson [Oke95], pages 113–118. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Kernighan:1990:ITD

[Ker90] Brian W. Kernighan. Issues and tradeoffs in document preparation systems. In Furuta [Fur90a], pages 1–16. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Keyhani:1995:ICR

[Key95] A. Keyhani. Innovation in cost recovery. In Okerson [Oke95], pages 57–66. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Kumar:1995:IIK

[KFA95] V. Kumar, R. Furuta, and R. B. Allen. Interactive interfaces for knowledge-rich domains. In Brown et al. [BBKF95], pages 235–246. ISSN 0894-3982.

King:1995:MMD

[Kin95] P. R. King. Modelling multimedia documents. In Brown et al. [BBKF95], pages 95–110. ISSN 0894-3982.

Klassen:1993:VWS

[Kla93] R. V. Klassen. Variable width splines: a possible font representation? In André et al. [AGS93], pages 183–194. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Kahn:1990:DHP

[KLLP90] Paul Kahn, Julie Launhardt, Krysztof Lenk, and Ronnie Peter. Design of hypermedia publications: Issues and solutions. In Furuta [Fur90a], pages 107–124. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Kilpelainen:1990:SDD

[KLMN90] Pekka Kilpeläinen, Greger Lindén, Heikki Mannila, and Erja Nikunen. A structured document database system. In Furuta [Fur90a], pages 139–151. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Kohen:1989:SEW

[Koh89] E. Kohen. A simple and efficient way to design middle resolution fonts. In André and Hersch [AH89], pages 22–33. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Koenig:1993:AEP

[KOOR93] J. A. Koenig, D. B. Orr, E. Ozil, and D. Rodgers. The AMS electronic publishing experiment: A new vision of the scientific journal (a summary). In Okerson [Oke93], pages 89–90. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Kuikka:1993:TSD

[KP93] E. Kuikka and M. Penttonen.

Transformation of structured documents with the use of grammar. In Hüser et al. [HMQ93], pages 373–384. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Kuikka:1995:FSD

[KS95] E. Kuikka and A. Salminen. Filtering structured documents in the SYNDOC environment. In Brown et al. [BBKF95], pages 181–194. ISSN 0894-3982.

Lamberton:1993:CT

[Lam93] D. Lamberton. The cost of thinking. In Anonymous [Ano93g], pages 98–100. ISSN 0004-9670.

Lamarca:1997:EPU

[Lam97] M. M. Lamarca. Electronic publishing in a university setting: The centralization vs. decentralization debate. In Logan and Gluck [LG97], pages 81– 94. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Lawrence:1995:VPF

[Law95] J. Lawrence. Virtual publication and the fair use concept. In Okerson [Oke95], pages 219–228. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Lie:1992:EBA

[LB92] Hakon Lie and Walter Bender. The electronic broadsheet: all the news that fit the display. In Vanoirbeek and Coray [VC92a], pages 199–209. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Lovegrove:1995:DAP

[LB95] W. S. Lovegrove and D. F. Brailsford. Document analysis of PDF files: Methods, results and implications. In Brown et al. [BBKF95], pages 207–220. ISSN 0894-3982.

Lebrave:1992:LAT

[Leb92] Jean-Louis Lebrave. L’hypertexte et l’avant-texte (invited paper). In Vanoirbeek and Coray [VC92a], pages 233–246. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Levy:1993:DRD

[Lev93] D. M. Levy. Document reuse and document systems. In Hüser et al. [HMQ93], pages 339–348. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Levy:1995:NNN

[Lev95] D. Levy. Naming the namable:

Names, versions, and document identity in a networked environment. In Okerson [Oke95], pages 153–160. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Lima:1990:ISD

[LG90] José Valdeni De Lima and Henri Galy. The integration of structured documents into DMBS. In Furuta [Fur90a], pages 153–168. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Logan:1997:EPA

[LG97] Elisabeth Logan and Myke Gluck, editors. Electronic publishing: applications and implications: contributed papers to the ASIS midyear meeting, Minneapolis, MN, May 1995. Information Today, Medford, NJ, USA, 1997. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Lewis:1995:PMT

[LK95] S. Lewis and T. Kelley. Project Muse: Tackling 40 journals. In Okerson [Oke95], pages 103–112. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

LeVan:1986:LDF

[LT86] H. Le Van and E. Terreni. A language to describe formatting directives for SGML documents. In Désarménien [Dés86], pages 98– 119. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Lucarella:1986:RMF

[Luc86] D. Lucarella. Retrieving mathematical formulae. In Désarménien [Dés86], pages 120–130. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Luckhurst:1997:WWP

[Luc97] G. Luckhurst. Who wants peer review anyway? In Secker [Sec97], pages 23–34. ISBN 0-85296-691-1. LCCN ????

Lyons:1995:ADO

[Lyo95] P. A. Lyons. Access to digital objects: A communications law perspective. In Okerson [Oke95], pages 213–218. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Morris:1991:RID

[MA91] Robert A. Morris and Jacques André, editors. Raster Imaging and Digital Topography II: Papers from the second RIDT meeting, held in Boston, Oct. 14–16, 1991. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Maag:1989:SIB

[Maa89] B. Maag. Shape investigation with bitmapped characters. In André and Hersch [AH89], pages 180–187. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

MacKay:1991:LPQ

[Mac91] Pierre A. MacKay. Looking at the pixels: Quality control for 300 dpi laser printer fonts, especially s. In Morris and André [MA91], pages 205– 215. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

MacLachlan:1997:LTR

[Mac97] L. MacLachlan. Libraries: Turmoil and renaissance. In Secker [Sec97], pages 35–40. ISBN 0-85296-691-1. LCCN ????

McQueen:1993:IPF

[MB93] C. D. McQueen and R. G. Beausoleil. Infinifont: a parametric font generation system. In André et al. [AGS93], pages 117–132. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Morris:1993:CEL

[MBM93] R. A. Morris, E. M. Blachman, and C. Meyer. A constraint-based editor for linguistic scholars. In Hüser et al. [HMQ93], pages 349– 360. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

McLean:1993:PSI

[McL93] N. McLean. Provision of scholarly information: The economics of supply and demand. In Anonymous [Ano93g], pages 94– 97. ISSN 0004-9670.

McMillan:1995:SCP

[McM95] G. McMillan. Scholarly communications project: Publishers and libraries. In Okerson [Oke95], pages 135–146. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

McNicol:1993:APF

[McN93] D. McNicol. An administrator’s perspective on funding the electronic library. In Anonymous [Ano93g], pages 108–111. ISSN 0004-9670.

Matylonek:1997:HEP

[MGLD97] J. C. Matylonek, J. L. Green, E. Liss, and A. Duncan. Hort-Base: An example of professional societies’ roles in electronic information systems. In Logan and Gluck [LG97], pages 95– 106. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Murata:1992:FHS

[MH92] Makoto Murata and Koichi Hayashi. Formatter hierarchy for structured documents. In Vanoirbeek and Coray [VC92a], pages 77–94. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Moehr:1993:C

[MH93] W. Moehr and C. Hueser. Colophon. In Hüser et al. [HMQ93], pages 519–520. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Midwinter:1997:SUW

[Mid97] J. Midwinter. Strategies for the user — what are computers good for? or why do we still have paper? In Secker [Sec97], pages 65– 74. ISBN 0-85296-691-1. LCCN ????

Macleod:1992:FDS

[MNBH92] Ian A. Macleod, Brent Nordin, David T. Barnard, and Doug Hamilton. A framework for developing SGML applications. In Vanoirbeek and Coray [VC92a], pages 53–63. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Mittelbach:1992:PQH

[MR92] Frank Mittelbach and Chris Rowley. The pursuit of quality: How can automated typesetting achieve the highest standards of craft typography? In Vanoirbeek and Coray [VC92a], pages 261– 273. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Moon:1990:CFT

[MS90] Y. S. Moon and T. Y. Shin. Chinese fonts and their digitization. In Furuta [Fur90a], pages 235– 248. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Munson:1995:NPL

[Mun95] E. V. Munson. A new presentation language for structured documents. In Brown et al.

[BBKF95], pages 125–138. ISSN 0894-3982.

Mitsa:1991:CEH

[MUP91] Theophano Mitsa, Robert Ulichney, and Kevin J. Parker. The construction and evaluation of halftone patterns with manipulated power spectra. In Morris and André [MA91], pages 90– 97. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Mylonas:1993:PP

[Myl93] E. Mylonas. The Perseus Project. In Okerson [Oke93], pages 81– 88. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Naiman:1991:CSN

[Nai91] Avi C. Naiman. CRT spatial non-linearities and luminance linearization. In Morris and André [MA91], pages 43–53. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Nanard:1989:AAM

[NN89] M. Nanard and J. Nanard. An anti-aliasing method for low resolution fonts based on font structure. In André and Hersch [AH89], pages 111–122. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Nanard:1989:DAF

[NNGP89] M. Nanard, J. Nanard, M. Gandara, and N. Porte. Declarative approach for font design by incremental learning. In André and Hersch [AH89], pages 71– 82. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Noordzij:1991:SS

[Noo91] Gerrit Noordzij. The shape of the stroke. In Morris and André [MA91], pages 34–42. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Oakley:1993:CEP

[Oak93] R. L. Oakley. Copyright and electronic publishing. In Okerson [Oke93], pages 91–100. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

ODonnell:1993:TPT

[O’D93] J. J. O’Donnell. Time present and time past; an introduction to the symposium. In Okerson [Oke93], pages 1–4. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Okerson:1993:SPE

[Oke93] Anne Okerson, editor. Scholarly publishing on the electronic networks: the new generation: visions and opportunities in notfor-profit publishing: proceedings of the second symposium, December 5–8, 1992, the Washington Vista Hotel, Washington, DC, USA. Office of Scientific and Academic Publishing, Association of Research Libraries, Washington, DC, USA, 1993. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Okerson:1995:FPP

[Oke95] Anne Okerson, editor. Filling the pipeline and paying the piper: proceedings of the fourth symposium, November 5–7, 1994, the Washington Vista Hotel, Washington, DC. Association of Research Libraries, Washington, DC 20036, USA, 1995. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Ou:1989:FGA

[OO89] C. Ou and Y. Ohno. Font generation algorithms for Kanji characters. In André and Hersch [AH89], pages 123–133. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Owens:1993:ETS

[Owe93] E. Owens. Electronic text and scholarly publishers: How and why? In Okerson [Oke93], pages 65–72. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Perelman:1995:KSA

[Per95] L. J. Perelman. KEYNOTE: Is school out? is academic publishing out? In Okerson [Oke95], pages 15–26. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Petersen:1986:TDF

[Pet86] H. W. Petersen. The TEX document factory in a university environment: process model, implementation steps, experiences. In Désarménien [Dés86], pages 136– 144. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Penny:1986:ITE

[PH86] P. Penny and J. L. Henriot. Integrating TEX in an EDDS with very high resolution capabilities. In Désarménien [Dés86], pages 131–135. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Pitti:1995:BFA

[Pit95] D. V. Pitti. The Berkeley Finding Aids Project; standards in navigation. In Okerson [Oke95], pages 161–166. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Porter:1992:DRS

[PR92] Gilbert B. Porter, III and Emil V.

Rainero. Document reconstruction: A system for recovering document structure from layout. In Vanoirbeek and Coray [VC92a], pages 127–141. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Quint:1990:TDE

[QNA90] Vincent Quint, Marc Nanard, and Jacques André. Towards document engineering. In Furuta [Fur90a], pages 17–29. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Quint:1986:GIE

[QVB86] V. Quint, I. Vatton, and H. Bedor. GRIF: An interactive environment for TEX. In Désarménien [Dés86], pages 145– 158. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Richy:1995:TSS

[RA95] H. Richy and J. Andre. Typographic sheets and structured documents. In Brown et al. [BBKF95], pages 81–94. ISSN 0894-3982.

Richy:1992:MSS

[RFP92] Hélène Richy, Patrice Frison, and Eric Picheral. Multilingual stringto-string correction in Grif, a structured editor. In Vanoirbeek and Coray [VC92a], pages 183– 198. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Rossiter:1990:SWD

[RH90] B. N. Rossiter and M. A. Heather. Strengths and weaknesses of database models for textual documents. In Furuta [Fur90a], pages 125–138. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Rostek:1993:WWS

[RMF93] L. Rostek, W. Moehr, and D. H. Fischer. Weaving a web: The structure and creation of an object network representing an electronic reference work. In Hüser et al. [HMQ93], pages 495–506. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Rodgers:1995:SSP

[Rod95] D. L. Rodgers. Scientific scholarly publishing: A draft proposal. In Okerson [Oke95], pages 179– 182. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Rohrich:1986:AMT

[Roh86] J. Rohrich. Abstract markup in TEX. In Désarménien [Dés86], pages 159–160. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Rosenberg:1991:LCM

[Ros91] Charles Rosenberg. A low complexity method of compressing Kanji font bitmaps. In Morris and André [MA91], pages 137– 146. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Rous:1993:EPF

[Rou93] B. Rous. Electronic publishing; A five-year plan. In Okerson [Oke93], pages 29–42. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Rubinsky:1993:ETD

[Rub93] Y. Rubinsky. Electronic texts the day after tomorrow. In Okerson [Oke93], pages 5–14. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Roisin:1993:MLP

[RV93] C. Roisin and I. Vatton. Merging logical and physical structures in documents. In Hüser et al. [HMQ93], pages 327–338. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Salton:1992:ADC

[SAB92] Gerard Salton, James Allan, and Chris Buckley. Automatic determination of content relationships in natural-language texts. In Vanoirbeek and Coray [VC92a], pages 165–181. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Sampson:1997:CWI

[Sam97] P. Sampson. Copyright — will it help or hinder. In Secker [Sec97], pages 51–58. ISBN 0-85296-691-1. LCCN ????

Scheermesser:1993:DHT

[SB93] T. Scheermesser and O. Bryngdahl. Digital halftoning with texture control. In André et al. [AGS93], pages 207–212. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Smith:1995:TSB

[SB95] P. N. Smith and D. F. Brailsford. Towards structured, block-based PDF. In Brown et al. [BBKF95], pages 153–166. ISSN 0894-3982.

Smith:1993:JPA

[SBEH93] P. N. Smith, D. F. Brailsford, D. R. Evans, and L. Harrison. Journal publishing with Acrobat: the CAJUN Project. In Hüser et al. [HMQ93], pages 481–494. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Seaman:1993:ETC

[Sea93] D. Seaman. The Electronic Text Center; on-line archive of electronic texts at the University of Virginia libraries. In Okerson [Oke93], pages 101–106. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Seaman:1995:CPS

[Sea95] D. Seaman. Campus publishing in standardized electronic formats — HTML and TEI. In Okerson [Oke95], pages 93–102. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Secker:1997:EPH

[Sec97] Philip Secker, editor. Electronic publishing: how will it work?: edited proceedings of a forum held on 13 May 1997, organised by the Institution of Electrical Engineers as part of the Engineering Council 20/20 Vision Programme, Communications theme. IEE, London, UK, 1997. ISBN 0-85296-691-1. LCCN ????

Seitter:1995:FSO

[Sei95] K. Seitter. Five societies: One journal project. In Okerson [Oke95], pages 147–152. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Sennhauser:1993:IRA

[Sen93] R. Sennhauser. Improving the recognition accuracy of text recognition systems using typographical constraints. In André et al. [AGS93], pages 273–282. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Suellow:1993:MFI

[SGBOP93] K. Suellow, I. Gabel-Becker, M. Ockenfeld, and W. Putz. MultiMedia forum: an interactive online journal. In Hüser et al. [HMQ93], pages 413–422. ISBN 0-471-95070-X. ISSN 0894-3982. LCCN ????

Stephen:1993:CEC

[SH93] T. Stephen and T. M. Harrison.

Comserve: An electronic community for communication scholars.

In Okerson [Oke93], pages 53– 58. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Shipp:1993:WAO

[Shi93] J. Shipp. The way ahead: Outcomes of the Changes in scholarly communication patterns conference, Canberra, 14–16 April 1993. In Anonymous [Ano93g], pages 84–88. ISSN 0004-9670.

Shaw:1996:EPS

[SM96] Dennis Shaw and Howard Moore, editors. Electronic publishing in science: proceedings of the Joint ICSU Press/Unesco Expert Conference: February 1996, Paris, France, volume 9(4) of Learned Publishing. Association of Learned and Professional Society Publishers and Unesco, Hatfield, Herts, UK, 1996. ISBN 0-930357-37-X. ISSN 0953-1513. LCCN ????

Story:1993:RIE

[SOFS93] G. A. Story, L. O’Gorman, D. Fox, and L. Schaper. The RightPages image-based electronic library for alerting and browsing (reprinted by permission of IEEE, Computer, September 1992). In Okerson [Oke93], pages 123–132. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Solomon:1993:SNJ

[Sol93] D. L. Solomon. Starting a new journal in statistics education. In Okerson [Oke93], pages 19– 28. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Southall:1986:DNT

[Sou86] Richard Southall. Designing a new typeface with METAFONT. In Désarménien [Dés86], pages 161–179. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Southall:1991:CDT

[Sou91] Richard Southall. Character description techniques in type manufacture. In Morris and André [MA91], pages 16–27. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Southall:1992:PRR

[Sou92] Richard Southall. Presentation rules and rules of composition in the formatting of complex text. In Vanoirbeek and Coray [VC92a], pages 275–290. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Southwick:1995:UGI

[Sou95] T. Southwick. The U.S. Government’s interest in copyright and fair use (Executive Summary of the Report of the NII Working Group on Intellectual Property Rights). In Okerson [Oke95], pages 189–194. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Spender:1993:IPR

[Spe93] L. Spender. Intellectual property rights in an electronic age. In Anonymous [Ano93g], pages 101– 103. ISSN 0004-9670.

Spitz:1990:RPM

[Spi90] A. Lawrence Spitz. Recognition processing for multilingual document. In Furuta [Fur90a], pages 193–205. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Spink:1997:UBR

[SS97] A. Spink and L. Schamber. Use of book reviews by scholars: Implications for electronic publishing. In Logan and Gluck [LG97], pages 113–130. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Stamm:1993:DRI

[Sta93a] B. Stamm. Dynamic regularisation of intelligent outline fonts. In André et al. [AGS93], pages 219– 230. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Stamm:1993:OEF

[Sta93b] B. Stamm. Object-orientation and extensibility in a font-scaler. In André et al. [AGS93], pages 159–170. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Semenzato:1992:ALT

[SW92] Luigi Semenzato and Edward Wang. Aleph: A language for typesetting. In Vanoirbeek and Coray [VC92a], pages 65–75. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Shepherd:1997:EN

[SWB97] M. A. Shepherd, C. R. Watters, and F. J. Burkowski. Electronic news. In Logan and Gluck [LG97], pages 67–80. ISBN 1-57387-036-6. LCCN Z286.E43L64 1997.

Tamari:1989:DHF

[Tam89] I. J. Tamari. Digitalisation of hebrew fonts, or: Some evolutional evaluations. In André and Hersch [AH89], pages 188–197. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Tamari:1991:DLR

[Tam91] Ittai Tamari. Decipherability, legibility and readability of modern Hebrew typefaces. In Morris and André [MA91], pages 128– 136. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Tattersall:1992:MCT

[TC92] Colin Tattersall and Andrew J.

Cole. Modelling the content of technical documentation. In Vanoirbeek and Coray [VC92a], pages 223–232. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Tavakoli:1992:NMF

[TR92] Nassrin Tavakoli and Alan Ray.

A new method for full-text retrieval. In Vanoirbeek and Coray [VC92a], pages 153–163. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Uchio:1989:MNA

[UHST89] F. Uchio, T. Higuchi, H. Sanada, and Y. Tezuka. A method of normalizing appearance size of brush-written Chinese characters. In André and Hersch [AH89], pages 144–153. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

vanVliet:1988:DMT

[van88] J. C. van Vliet, editor. Document Manipulation and Typography. Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Nice (France), April 20–22, 1988. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Vanoirbeek:1992:FST

[Van92] Christine Vanoirbeek. Formatting structured tables (invited paper). In Vanoirbeek and Coray [VC92a], pages 291–309. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

VanSteenberg:1993:SEE

[Van93] M. E. Van Steenberg. STE-LAR: An experiment in the electronic distribution of astronomical literature. In Okerson [Oke93], pages 43–52. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Vanoirbeek:1992:EPE

[VC92a] C. Vanoirbeek and G. Coray, editors. EP92—Proceedings of Electronic Publishing, ’92, International Conference on Electronic Publishing, Document Manipulation, and Typography, Swiss Federal Institute of Technology, Lausanne, Switzerland, April 7-10, 1992. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Vanoirbeek:1992:C

[VC92b] Christine Vanoirbeek and Giovanni Coray. Colophon. In Vanoirbeek and Coray [VC92a], pages 311–312. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Vercoustre:1990:SEH

[Ver90] Anne-Marie Vercoustre. Structured editing—hypertext approach: Cooperation and complementarity. In Furuta [Fur90a], pages 65–78. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

VanOssenbruggen:1995:WAS

[VES95] J. Van Ossenbruggen, A. Eliens, and B. Schoenhage. Web applications and SGML. In Brown et al. [BBKF95], pages 51–62. ISSN 0894-3982.

Varian:1995:SFA

[VM95] H. R. Varian and J. K. MacKie Mason. Some FAQs about usage-based pricing. In Okerson [Oke95], pages 77–86. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

vanBlokland:1991:DAL

[vv91] Erik van Blokland and Just van Rossum. Different approaches to lively outlines. In Morris and André [MA91], pages 28– 33. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Wolfman:1990:FLT

[WB90] Anthony P. Wolfman and Daniel M. Berry. flo—A language for typesetting flowcharts. In Furuta [Fur90a], pages 93–106. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Wonneberger:1986:VVT

[Won86] R. Wonneberger. “verheissung und versprechen”, a third generation approach to theological typesetting. In Désarménien [Dés86], pages 180–198. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Wuthrich:1989:FDA

[WSG89] C. A. Wüthrich, P. Stucki, and A. Ghezal. A frequency domain analysis of square- and hexagonal- grid-based images. In André and Hersch [AH89], pages 261–270. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Wang:1995:XTE

[WW95] X. Wang and D. Wood. Xtable — A tabular editor and formatter. In Brown et al. [BBKF95], pages 167–180. ISSN 0894-3982.

Zapf:1993:AMH

[Zap93] H. Zapf. About micro-typography and the hz-program. In André et al. [AGS93], pages 283–288. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Zeggel:1993:EDA

[ZB93] T. Zeggel and O. Bryngdahl.

Error diffusion on an adaptive raster. In André et al. [AGS93], pages 213–218. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.

Zhang:1995:PLC

[ZD95] Q. Zhang and J. M. Danskin. A pattern-based lossy compression scheme for document. In Brown et al. [BBKF95], pages 221–234. ISSN 0894-3982.

Zheng:1991:SRM

[Zhe91]  Youlu Zheng.  Subdivision-

remainder method for converting spatial resolution to intensity resolution. In Morris and André [MA91], pages 81–89. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Zheng:1992:DCE

[Zhe92] Min Zheng. A document composition environment for multilanguage processing. In Vanoirbeek and Coray [VC92a], pages 43–52. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Zramdini:1993:OFR

[ZI93] A. Zramdini and R. Ingold. Optical font recognition from projection profiles. In André et al. [AGS93], pages 249–260. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN Z250.7.